# Sim-DETR: Unlock DETR for Temporal Sentence Grounding

### v1

@@ Abstract
Temporal sentence grounding aims to identify exact moments in a video that correspond to a given textual query, typically addressed with detection transformer (DETR) solutions. However, we find that typical strategies designed to enhance DETR do not improve, and may even degrade, its performance in this task. We systematically analyze and identify the root causes of this abnormal behavior: (1) conflicts between queries from similar target moments and (2) internal query conflicts due to the tension between global semantics and local localization. Building on these insights, we propose a simple yet powerful baseline, Sim-DETR, which extends the standard DETR with two minor modifications in the decoder layers: (1) constraining self-attention between queries based on their semantic and positional overlap and (2) adding query-to-frame alignment to bridge the global and local contexts. Experiments demonstrate that Sim-DETR unlocks the full potential of DETR for temporal sentence grounding, offering a strong baseline for future research. Code is available at https://github.com/nagara214/Sim-DETR.
@@ 1 Introduction
Video has become a dominant media on the internet, with short-form content rapidly expanding and achieving exponential growth in reach over recent years [18, 24, 19, 20]. Instead of passively watching entire videos, users now prefer to target specific segments of interest, with natural language descriptions serving as an intuitive and flexible way to convey intent [61, 30, 68]. This amplifies interest in the research topic of temporal sentence grounding [71, 57, 40, 22, 32, 47, 27, 46, 25, 81, 26], which aims to locate one or more semantically aligned segments within an untrimmed video according to a given natural language sentence, as shown in Figure 1a.
Early methods either align sentences with predefined temporal proposals for selection [44, 70] or directly predict moment spans through cross-modal interactions between language and frames [36, 37, 72]. Recent advances center on integrating the Detection Transformer (DETR) [2] for object detection into the temporal sentence grounding, leveraging its query-based proposal detection framework to eliminate hand-crafted proposal components and deliver superior grounding performance with high efficiency [22, 25, 57, 46, 47, 26]. They leverage moment queries—often several times more numerous than ground truth segments—in the multi-layer decoder, which are initialized randomly [46, 57, 26], generated from sentences [47, 22], or extracted as event units from the video [25], to locate and search the most suitable matches for target segments. A one-to-one label assignment is applied in the training objective to facilitate non-redundant predictions for each ground-truth segment. The overall architecture of the DETR-based temporal sentence grounding framework is shown in Figure 6a.
@@ 2 Related Work
Temporal Sentence Grounding (TSG) aims to identify specific video moments that align with given language expressions. As computer vision advances across neural architectures, generation, detection, segmentation, and multimodal tasks [4, 21, 55, 73, 33, 8, 54, 82, 53], temporal understanding requires task-specific optimizations. Early methods fall into proposal-free and proposal-based categories. Proposal-free methods use end-to-end frameworks by directly regressing boundary coordinates [43, 6, 76] or predicting frame-level boundary likelihood [16, 78]. Proposal-based approaches densely sample candidate segments via sliding windows [12, 1, 38, 80, 15, 23], temporal anchors [32, 41, 5, 77, 75, 81], or multimodal feature similarity [69, 7, 66, 35, 34]. Recently, DETR has shown exceptional effectiveness in detection tasks. [27, 65] pioneered DETR for TSG. Subsequent studies enhance cross-modal representations [47, 57, 67, 46], improve decoder queries with semantic/spatial information [22, 25, 58], or explore joint training with other temporal tasks [57, 71].
Detection Transformers (DETR) [2] is a pioneering end-to-end framework for object detection and segmentation that employs transformers to make direct set-based predictions [14, 85, 52, 55, 54, 61, 60]. However, DETR has notable limitations, especially its high computational complexity and slow convergence, impacting its practical efficiency. To address these issues, several derivative models have emerged. Some methods [84, 74, 63, 51] improve computational efficiency by reducing redundant calculations within the DETR architecture. Other models [13, 59, 9] refine the attention mechanism for more efficient information processing and resource utilization. Additionally [45, 64, 39] enhance convergence by embedding spatial information directly into the query design, which accelerates object localization. Recent models [28, 79] tackle optimization challenges in DETR’s bipartite matching process by integrating denoising strategies during training, improving both convergence and performance. Despite recent advancements in enhancing DETR for object detection, there is a lack of systematic studies addressing the unique challenges of applying DETR to TSG and the potential difficulties in query optimization.
@@ 3 Probing DETR’s Inherent Behavior for TSG
This study begins with the observation that enhancements effective for object detection in DETR [2] do not apply to temporal sentence grounding (TSG), as discussed in Sec 3.1. To explore the reasons behind DETR’s failure to enhance, this section presents preliminary studies to reveal the behavior of queries in the DETR decoder, as queries are central to predicting proposals. Specifically, we examine the similarity between queries and outputs in Sec 3.2, and the attention between queries and inputs (i.e., frames) in Sec 3.3. These help reveal the underlying cause, and our modification addressing it not only eliminates the abnormal phenomena but also has a positive side effect (Sec 3.4).
Experimental Settings. Unless otherwise specified, our default experimental setup involves conducting statistical analyses on the full validation set of QVHighlights [27]. CG-DETR [46] and TaskWeave [71], which serve as baselines for DETR-based TSG due to their strong performance while retaining the core vanilla DETR mechanism, are used with their standard configurations.
@@ 3.1 Abnormal Phenomena
Result & Discussion. As shown in Figure 2, increasing the number of queries to 15 and 20 led to progressively larger performance declines, with a drop of over -2% with TaskWeave [71] at 20 queries. Similarly, increasing the number of decoder layers caused a drop from -1% to -2.5%. In contrast, our Sim-DETR mitigates this phenomenon and achieves a slight improvement, demonstrating its robustness. This observation motivates an in-depth exploration of decoder layer, particularly the role of queries (see Sec 3.2).
@@ 4 Method
We present a simple yet effective TSG baseline that resolves query conflicts in existing DETR-based TSG frameworks [27, 47, 22, 46, 57, 67, 71, 25] through two minor yet key modifications. First, we introduce the standard DETR-based [2] TSG architecture (Sec 4.1). Sec 4.2 introduces our first modification, which adjusts the self-attention mechanism between queries in the decoder to avoid conflicts among them, based on the observations presented in Sec 3.2. Next, Sec 4.3 presents another improvement that addresses the conflict between global matching and local localization within the query (discussed in Sec 3.3) by introducing query-to-frame matching loss, which serves as the bridge between the global and local perspectives.
@@ 4.1 Standard DETR-based TSG Architecture
Feature Extraction. Previous DETR-based works [47, 57, 71, 25] employ CLIP [49] and SlowFast [10] to extract video features, represented as $\mathcal{T}\in\mathbb{R}^{N\times C}$, where $N$ and $C$ stands for the number of frames and the hidden dimension, respectively. For the language expression, typically a sentence with $L$ tokens, the CLIP text encoder is used to obtain token-level features, resulting in $\mathcal{W}\in\mathbb{R}^{L\times C}$.
Multimodal Encoder. After feature extraction, DETR-based TSG methods [47, 22, 46, 57] employ a multimodal encoder to fuse these video and language features, producing multimodal video features, denoted as $\hat{\mathcal{T}}\in\mathbb{R}^{N\times C}$. The key component of the multimodal encoder is a cross-attention mechanism [62], where video features $\mathcal{T}$ act as queries, while word features $\mathcal{W}$ serve as keys and values.
@@ 4.2 Query Grouping and Ranking
Meanwhile, to enhance cross-layer matching consistency, we draw inspiration from EASE-DETR [14]’s query ranking strategy in object detection. By ranking queries within a segment at each layer, the leading query is prioritized for segment matching in subsequent layers. However, directly ranking queries by prediction scores, as in EASE-DETR, proves ineffective (see Appendix D). As analyzed in Sec. 3.3, high prediction/confidence scores do not guarantee precise predictions and may correspond to poor localization. Instead, we introduce an IoU head to assess localization accuracy and rank queries based on both factors, as follows:
@@
Finally, we incorporate query grouping within segments and intra-segment ranking to infer relative relationships, then reshape query-to-query self-attention following [14] to mitigate conflicts, as follows:
@@ 4.3 Global-Local Bridging
M-DETR [27]NeurIPS’21 52.89 33.02 54.82 29.40 30.73 53.94 34.84 - - 32.20
UMT [40]CVPR’22 56.23 41.18 53.83 37.01 36.12 60.26 44.26 - - 38.59
QD-DETR [47]CVPR’23 62.40 44.98 62.52 39.88 39.86 62.68 46.66 62.23 41.82 41.22
UniVTG [32]ICCV’23 58.86 40.86 57.60 35.59 35.47 59.74 - - - 36.13
EaTR [22]ICCV’23 - - - - - 61.36 45.79 61.86 41.91 41.74
MomentDiff [29]NeurIPS’23 57.42 39.66 54.02 35.73 35.95 - - - - -
TR-DETR [57]AAAI’24 64.66 48.96 63.98 43.73 42.62 67.10 51.48 66.27 46.42 45.09
UVCOM [67]CVPR’24 63.55 47.47 63.37 42.67 43.18 65.10 51.81 - - 45.79
TaskWeave [71]CVPR’24 - - - - - 64.26 50.06 65.39 46.47 45.38
BAM-DETR [25]ECCV‘24 62.71 48.64 64.57 46.33 45.36 65.10 51.61 65.41 48.56 47.61
CG-DETR [46]Arxiv‘24 65.43 48.38 64.51 42.77 42.86 67.35 52.06 65.57 45.73 44.93
SpikeMba [31]Arxiv‘24 64.13 49.42 - 43.67 43.79 65.32 51.33 - 44.96 44.84
@@
2D-TAN [81] 40.01 27.99 12.92 27.22 58.76 46.02 27.50 41.25
M-DETR [27] 37.97 24.67 11.97 25.49 65.83 52.07 30.59 45.54
MomentDiff [29] 44.78 33.68 - - - 55.57 32.42 -
QD-DETR [47] - - - - - 57.31 32.55 -
UniVTG [32] 51.44 34.97 17.35 33.60 70.81 58.01 35.65 50.10
CG-DETR [46] 52.23 39.61 22.23 36.48 70.43 58.44 36.34 50.13
UVCOM [67] - 36.39 23.32 - - 59.25 36.64 -
SpikeMba [31] 51.98 39.34 22.83 35.81 71.24 59.65 36.12 51.74
@@ 4.4 Overall Loss
The overall loss is: $\mathcal{L}=\mathcal{L}_{\text{MD}}+\lambda_{\text{bridge}}\mathcal{L}_{\text{bridge}}+\lambda_{\text{iou}}\mathcal{L}_{\text{iou}},$ where $\mathcal{L}_{\text{MD}}$ includes the L1, gIoU, classification, and saliency losses, as in Moment DETR [26]. $\mathcal{L}_{\text{iou}}$ is the loss for the IoU head introduced in Line-406, designed to assist in ranking queries using $\mathcal{P}^{\text{IoU}}_{i}$ and $\mathcal{P}^{\text{IoU}}_{j}$ in Eq 3. The IoU head implementation follows [25, 58]. Ablations for hyperparameters $\lambda_{\text{bridge}}$ and $\lambda_{\text{iou}}$ are detailed in Appendix E.
@@ 5.1 Datasets and Implementation Details
Datasets. Following [27, 47, 32, 67, 31], we conduct comprehensive evaluation experiments across three datasets: QVHighlights [27], Charades-STA [12], and TACoS [50]. The QVHighlights dataset contains 10,310 text expressions and 18,367 distinct events, with most expressions corresponding to multiple events. Charades-STA is derived from a collection of 9,848 indoor scene videos, encompassing 16,128 events.TACoS includes only 273 videos, with an average length approaching five minutes per video.
@@
Implementation Details. We set the input video resolution to $\times224$ and represent each video by concatenating the CLIP [49] [CLS] tokens with SlowFast [10] features. For the Charades-STA and TACoS datasets, we additionally constructed two experiments using I3D [3] and VGG [56] as visual encoders, with Glove [48] embeddings for text features. The model is trained for 200 epochs on a single NVIDIA A40 GPU, using the AdamW [42] optimizer with an initial learning rate of 1e-4. Note that, for a fair comparison, the total number of layers in our Sim-DETR is set to 6, the same as recent works [46, 22, 67].
@@
VGG [56] 2D-TAN [81] 40.94 22.85
FVMR [11] 42.36 24.14
SSRN [83] 46.72 27.98
UMT [40] 48.31 29.25
MomentDiff [29] 51.94 28.25
QD-DETR [47] 52.77 31.13
TR-DETR [57] 53.47 30.81
CG-DETR [46] 55.22 34.19
@@
I3D [3] MAN [77] 46.53 22.72
VSLNet [78] 47.31 30.19
QD-DETR [47] 50.67 31.02
TR-DETR [57] 55.51 33.66
TaskWeave [71] 53.36 31.40

### v2

@@ Abstract
Temporal sentence grounding aims to identify exact moments in a video that correspond to a given textual query, typically addressed with detection transformer (DETR) solutions. However, we find that typical strategies designed to enhance DETR do not improve, and may even degrade, its performance in this task. We systematically analyze and identify the root causes of this abnormal behavior: (1) conflicts between queries from similar target moments and (2) internal query conflicts due to the tension between global semantics and local localization. Building on these insights, we propose a simple yet powerful baseline, Sim-DETR, which extends the standard DETR with two minor modifications in the decoder layers: (1) constraining self-attention between queries based on their semantic and positional overlap and (2) adding query-to-frame alignment to bridge the global and local contexts. Experiments demonstrate that Sim-DETR unlocks the full potential of DETR for temporal sentence grounding, offering a strong baseline for future research. Code is available at https://github.com/SooLab/Sim-DETR.
@@ 1 Introduction
Video has become a dominant media on the internet, with short-form content rapidly expanding and achieving exponential growth in reach over recent years [han2020self, kim2021relational, huang2018makes, huang2023free38]. Instead of passively watching entire videos, users now prefer to target specific segments of interest, with natural language descriptions serving as an intuitive and flexible way to convey intent [tang2023temporal71, li2023progressive, xie2023ra]. This amplifies interest in the research topic of temporal sentence grounding [yang2024taskweave, sun2024trdetr, liu2022umt, jang2023eatr, lin2023univtg, moon2023qd-detr, lei2021qvhighlights, moon2023cgdetr, lee2025bamdetr, zhang2020tan, lei2021detecting], which aims to locate one or more semantically aligned segments within an untrimmed video according to a given natural language sentence, as shown in Figure 1a.
Early methods either align sentences with predefined temporal proposals for selection [mao2020generation, yadav2020unsupervised] or directly predict moment spans through cross-modal interactions between language and frames [liu2022memory, liu2022reducing, yang2022entity]. Recent advances center on integrating the Detection Transformer (DETR) [carion2020detr] for object detection into the temporal sentence grounding, leveraging its query-based proposal detection framework to eliminate hand-crafted proposal components and deliver superior grounding performance with high efficiency [jang2023eatr, lee2025bamdetr, sun2024trdetr, moon2023cgdetr, moon2023qd-detr, lei2021detecting]. They leverage moment queries—often several times more numerous than ground truth segments—in the multi-layer decoder, which are initialized randomly [moon2023cgdetr, sun2024trdetr, lei2021detecting], generated from sentences [moon2023qd-detr, jang2023eatr], or extracted as event units from the video [lee2025bamdetr], to locate and search the most suitable matches for target segments. A one-to-one label assignment is applied in the training objective to facilitate non-redundant predictions for each ground-truth segment. The overall architecture of the DETR-based temporal sentence grounding framework is shown in Figure 6a.
@@ 2 Related Work
Temporal Sentence Grounding (TSG) aims to identify specific video moments that align with given language expressions. As computer vision advances across neural architectures, generation, detection, segmentation, and multimodal tasks [chen2024survey12, huang2025mvtokenflow39, shi2024plain66, yang2021bottom81, lin2021structured46, dai2024curriculum20, shi2024part2object64, zheng2023ddcot88, shi2023logoprompt63], temporal understanding requires task-specific optimizations. Early methods fall into proposal-free and proposal-based categories. Proposal-free methods use end-to-end frameworks by directly regressing boundary coordinates [lu2019debug, chen2020gdp, zeng2020drn] or predicting frame-level boundary likelihood [ghosh2019excl, zhang2020vslnet]. Proposal-based approaches densely sample candidate segments via sliding windows [gao2017ctrl, anne2017mcn, liu2018role, zhang2019tcmn, ge2019mac, jiang2019slta], temporal anchors [lin2023univtg, liu2024tuning, chen2018tgn, zhang2019man, yuan2019scdm, zhang2020tan], or multimodal feature similarity [xu2019qspn, chen2019sap, xiao2021bpnet, liu2021apgn, liu2022slp]. Recently, DETR has shown exceptional effectiveness in detection tasks. [lei2021qvhighlights, woo2022explore] pioneered DETR for TSG. Subsequent studies enhance cross-modal representations [moon2023qd-detr, sun2024trdetr, xiao2024uvcom, moon2023cgdetr], improve decoder queries with semantic/spatial information [jang2023eatr, lee2025bamdetr, sun2024rgtr], or explore joint training with other temporal tasks [sun2024trdetr, yang2024taskweave].
Detection Transformers (DETR) [carion2020detr] is a pioneering end-to-end framework for object detection and segmentation that employs transformers to make direct set-based predictions [ease-detr, zhu2025rethinking, shi2023edadet62, shi2024plain66, shi2024part2object64, tang2023temporal71, tang2023contrastive70]. However, DETR has notable limitations, especially its high computational complexity and slow convergence, impacting its practical efficiency. To address these issues, several derivative models have emerged. Some methods [zhu2020deformable, yao2021efficient, wang2021pnp, roh2021sparse] improve computational efficiency by reducing redundant calculations within the DETR architecture. Other models [gao2021smca, sun2021tsp, dai2021dynamic] refine the attention mechanism for more efficient information processing and resource utilization. Additionally [meng2021conditional, wang2022anchor, liu2022dab] enhance convergence by embedding spatial information directly into the query design, which accelerates object localization. Recent models [li2022dn, zhang2022dino] tackle optimization challenges in DETR’s bipartite matching process by integrating denoising strategies during training, improving both convergence and performance. Despite recent advancements in enhancing DETR for object detection, there is a lack of systematic studies addressing the unique challenges of applying DETR to TSG and the potential difficulties in query optimization.
@@ 3 Probing DETR’s Inherent Behavior for TSG
This study begins with the observation that enhancements effective for object detection in DETR [carion2020detr] do not apply to temporal sentence grounding (TSG), as discussed in Sec 3.1. To explore the reasons behind DETR’s failure to enhance, this section presents preliminary studies to reveal the behavior of queries in the DETR decoder, as queries are central to predicting proposals. Specifically, we examine the similarity between queries and outputs in Sec 3.2, and the attention between queries and inputs (i.e., frames) in Sec 3.3. These help reveal the underlying cause, and our modification addressing it not only eliminates the abnormal phenomena but also has a positive side effect (Sec 3.4).
Experimental Settings. Unless otherwise specified, our default experimental setup involves conducting statistical analyses on the full validation set of QVHighlights [lei2021qvhighlights]. CG-DETR [moon2023cgdetr] and TaskWeave [yang2024taskweave], which serve as baselines for DETR-based TSG due to their strong performance while retaining the core vanilla DETR mechanism, are used with their standard configurations.
@@ 3.1 Abnormal Phenomena
Result & Discussion. As shown in Figure 2, increasing the number of queries to 15 and 20 led to progressively larger performance declines, with a drop of over -2% with TaskWeave [yang2024taskweave] at 20 queries. Similarly, increasing the number of decoder layers caused a drop from -1% to -2.5%. In contrast, our Sim-DETR mitigates this phenomenon and achieves a slight improvement, demonstrating its robustness. This observation motivates an in-depth exploration of decoder layer, particularly the role of queries (see Sec 3.2).
@@ 4 Method
We present a simple yet effective TSG baseline that resolves query conflicts in existing DETR-based TSG frameworks [lei2021qvhighlights, moon2023qd-detr, jang2023eatr, moon2023cgdetr, sun2024trdetr, xiao2024uvcom, yang2024taskweave, lee2025bamdetr] through two minor yet key modifications. First, we introduce the standard DETR-based [carion2020detr] TSG architecture (Sec 4.1). Sec 4.2 introduces our first modification, which adjusts the self-attention mechanism between queries in the decoder to avoid conflicts among them, based on the observations presented in Sec 3.2. Next, Sec 4.3 presents another improvement that addresses the conflict between global matching and local localization within the query (discussed in Sec 3.3) by introducing query-to-frame matching loss, which serves as the bridge between the global and local perspectives.
@@ 4.1 Standard DETR-based TSG Architecture
Feature Extraction. Previous DETR-based works [moon2023qd-detr, sun2024trdetr, yang2024taskweave, lee2025bamdetr] employ CLIP [radford2021clip] and SlowFast [feichtenhofer2019slowfast] to extract video features, represented as $\mathcal{T}\in\mathbb{R}^{N\times C}$, where $N$ and $C$ stands for the number of frames and the hidden dimension, respectively. For the language expression, typically a sentence with $L$ tokens, the CLIP text encoder is used to obtain token-level features, resulting in $\mathcal{W}\in\mathbb{R}^{L\times C}$.
Multimodal Encoder. After feature extraction, DETR-based TSG methods [moon2023qd-detr, jang2023eatr, moon2023cgdetr, sun2024trdetr] employ a multimodal encoder to fuse these video and language features, producing multimodal video features, denoted as $\hat{\mathcal{T}}\in\mathbb{R}^{N\times C}$. The key component of the multimodal encoder is a cross-attention mechanism [vaswani2017transformer], where video features $\mathcal{T}$ act as queries, while word features $\mathcal{W}$ serve as keys and values.
@@ 4.2 Query Grouping and Ranking
Meanwhile, to enhance cross-layer matching consistency, we draw inspiration from EASE-DETR [ease-detr]’s query ranking strategy in object detection. By ranking queries within a segment at each layer, the leading query is prioritized for segment matching in subsequent layers. However, directly ranking queries by prediction scores, as in EASE-DETR, proves ineffective (see Appendix D). As analyzed in Sec. 3.3, high prediction/confidence scores do not guarantee precise predictions and may correspond to poor localization. Instead, we introduce an IoU head to assess localization accuracy and rank queries based on both factors, as follows:
@@
Finally, we incorporate query grouping within segments and intra-segment ranking to infer relative relationships, then reshape query-to-query self-attention following [ease-detr] to mitigate conflicts, as follows:
@@ 4.3 Global-Local Bridging
M-DETR [lei2021qvhighlights]NeurIPS’21 52.89 33.02 54.82 29.40 30.73 53.94 34.84 - - 32.20
UMT [liu2022umt]CVPR’22 56.23 41.18 53.83 37.01 36.12 60.26 44.26 - - 38.59
QD-DETR [moon2023qd-detr]CVPR’23 62.40 44.98 62.52 39.88 39.86 62.68 46.66 62.23 41.82 41.22
UniVTG [lin2023univtg]ICCV’23 58.86 40.86 57.60 35.59 35.47 59.74 - - - 36.13
EaTR [jang2023eatr]ICCV’23 - - - - - 61.36 45.79 61.86 41.91 41.74
MomentDiff [li2023momentdiff]NeurIPS’23 57.42 39.66 54.02 35.73 35.95 - - - - -
TR-DETR [sun2024trdetr]AAAI’24 64.66 48.96 63.98 43.73 42.62 67.10 51.48 66.27 46.42 45.09
UVCOM [xiao2024uvcom]CVPR’24 63.55 47.47 63.37 42.67 43.18 65.10 51.81 - - 45.79
TaskWeave [yang2024taskweave]CVPR’24 - - - - - 64.26 50.06 65.39 46.47 45.38
BAM-DETR [lee2025bamdetr]ECCV‘24 62.71 48.64 64.57 46.33 45.36 65.10 51.61 65.41 48.56 47.61
CG-DETR [moon2023cgdetr]Arxiv‘24 65.43 48.38 64.51 42.77 42.86 67.35 52.06 65.57 45.73 44.93
SpikeMba [li2024spikemba]Arxiv‘24 64.13 49.42 - 43.67 43.79 65.32 51.33 - 44.96 44.84
@@
2D-TAN [zhang2020tan] 40.01 27.99 12.92 27.22 58.76 46.02 27.50 41.25
M-DETR [lei2021qvhighlights] 37.97 24.67 11.97 25.49 65.83 52.07 30.59 45.54
MomentDiff [li2023momentdiff] 44.78 33.68 - - - 55.57 32.42 -
QD-DETR [moon2023qd-detr] - - - - - 57.31 32.55 -
UniVTG [lin2023univtg] 51.44 34.97 17.35 33.60 70.81 58.01 35.65 50.10
CG-DETR [moon2023cgdetr] 52.23 39.61 22.23 36.48 70.43 58.44 36.34 50.13
UVCOM [xiao2024uvcom] - 36.39 23.32 - - 59.25 36.64 -
SpikeMba [li2024spikemba] 51.98 39.34 22.83 35.81 71.24 59.65 36.12 51.74
@@ 4.4 Overall Loss
The overall loss is: $\mathcal{L}=\mathcal{L}_{\text{MD}}+\lambda_{\text{bridge}}\mathcal{L}_{\text{bridge}}+\lambda_{\text{iou}}\mathcal{L}_{\text{iou}},$ where $\mathcal{L}_{\text{MD}}$ includes the L1, gIoU, classification, and saliency losses, as in Moment DETR [lei2021detecting]. $\mathcal{L}_{\text{iou}}$ is the loss for the IoU head introduced in Line-406, designed to assist in ranking queries using $\mathcal{P}^{\text{IoU}}_{i}$ and $\mathcal{P}^{\text{IoU}}_{j}$ in Eq 3. The IoU head implementation follows [lee2025bamdetr, sun2024rgtr]. Ablations for hyperparameters $\lambda_{\text{bridge}}$ and $\lambda_{\text{iou}}$ are detailed in Appendix E.
@@ 5.1 Datasets and Implementation Details
Datasets. Following [lei2021qvhighlights, moon2023qd-detr, lin2023univtg, xiao2024uvcom, li2024spikemba], we conduct comprehensive evaluation experiments across three datasets: QVHighlights [lei2021qvhighlights], Charades-STA [gao2017ctrl], and TACoS [regneri2013tacos]. The QVHighlights dataset contains 10,310 text expressions and 18,367 distinct events, with most expressions corresponding to multiple events. Charades-STA is derived from a collection of 9,848 indoor scene videos, encompassing 16,128 events.TACoS includes only 273 videos, with an average length approaching five minutes per video.
@@
Implementation Details. We set the input video resolution to $\times224$ and represent each video by concatenating the CLIP [radford2021clip] [CLS] tokens with SlowFast [feichtenhofer2019slowfast] features. For the Charades-STA and TACoS datasets, we additionally constructed two experiments using I3D [carreira2017i3d] and VGG [simonyan2014vgg] as visual encoders, with Glove [pennington2014glove] embeddings for text features. The model is trained for 200 epochs on a single NVIDIA A40 GPU, using the AdamW [loshchilov2017adamw] optimizer with an initial learning rate of 1e-4. Note that, for a fair comparison, the total number of layers in our Sim-DETR is set to 6, the same as recent works [moon2023cgdetr, jang2023eatr, xiao2024uvcom].
@@
VGG [simonyan2014vgg] 2D-TAN [zhang2020tan] 40.94 22.85
FVMR [gao2021fvmr] 42.36 24.14
SSRN [zhu2023ssrn] 46.72 27.98
UMT [liu2022umt] 48.31 29.25
MomentDiff [li2023momentdiff] 51.94 28.25
QD-DETR [moon2023qd-detr] 52.77 31.13
TR-DETR [sun2024trdetr] 53.47 30.81
CG-DETR [moon2023cgdetr] 55.22 34.19
@@
I3D [carreira2017i3d] MAN [zhang2019man] 46.53 22.72
VSLNet [zhang2020vslnet] 47.31 30.19
QD-DETR [moon2023qd-detr] 50.67 31.02
TR-DETR [sun2024trdetr] 55.51 33.66
TaskWeave [yang2024taskweave] 53.36 31.40